OpenStudy (anonymous):

hey I need help to factorise, can someone help me with this equation, x^2-16+16?

OpenStudy (campbell_st):

ok so its \[x^2 -16 + 16 = x^2\] is the expression correctly typed?

OpenStudy (anonymous):

yeah that's what I meant :)

OpenStudy (campbell_st):

well you end up with x^2 = x* x

OpenStudy (lgbasallote):

i think he meant \[x^2 - 16x + 16\]

OpenStudy (campbell_st):

lol... ok... good luck

OpenStudy (anonymous):

\[x=(-b \pm \sqrt{b^2-4ab})/2a\]

OpenStudy (anonymous):

have you got it?

OpenStudy (anonymous):

yeah sorry, I just meant\[x^{2}\]-16x+16,not x^2-16x+16=x^2

OpenStudy (anonymous):

i just need to do that equation and then solve for x, i'm still a bit stuck

OpenStudy (anonymous):

you have to find x which can be found by using quadratic formula which I just have written it can not factorize directly

OpenStudy (anonymous):

okay, thanks for your help